OpenStudy (anonymous):

Caitlin had $402 in her bank account. She withdrew $15 each week to pay for a swimming lesson. She now has $237. a. Write an equation that can be used to find the number of swimming lessons that she paid for. b. How many swimming lessons did she pay for? c. At the time she had $237, the cost of a lesson rose to $19. How many lessons can she pay for with her remaining $237?

OpenStudy (anonymous):

237 = 402 - 15x when x is the number of weeks she withdrew 15 dollars

OpenStudy (anonymous):

then for c, 0 = 237 - 19x if you get x to be a fraction, round down to the nearest whole number

OpenStudy (anonymous):

but that would mean that a would be like this 237=402-15x -402 -402 -165=15x -165/15=x -11=x

OpenStudy (anonymous):

on your third line it should be -165 = -15x

OpenStudy (anonymous):

Ohh i see now! and can you just explain c for me a little bit more. all i did was divide -237/-19 and got 12.47. then just round down so it would be 12?

OpenStudy (anonymous):

Yes, because thats the number of whole lessons she can pay for, she can pay for almost an additional half of a lesson, but I don't think thats what they are looking for.